Suppose you collect the following macroeconomic data for Tuttleland (T$ is the symbol for Tuttlebucks):

Year Money Supply Proportion of Nominal Income Held as Money Real GDP
2008 T$300 Billion 0.20 T$1,500 Billion
2018 T$2,000 Billion 0.25 T$2,975 Billion

Calculate the rate of growth in the money supply, proportion of nominal income held as money, and real GDP. Also, using the rate of growth version of the quantity equation, calculate the rate of growth in nominal GDP, the rate of growth in prices (inflation). Finally, calculate the average annual inflation rate.
Answer #1
  • Proportion of Nominal Money held as money = Money Supply * Proportion of nominal money held

Thus Proportion of Nominal Money held as money in 2008 = 0.20 * 300 = 60 Billion

Proportion of Nominal Money held as money in 2018 = 0.25 * 2000 = 500 Billion

  • Money supply in 2008 = 300

Money supply in 2018 = 2000

Growth rate in money supply = [(2000 - 300) / 300] * 100 = 566.67%

  • Real GDP in 2008 = 1500

Real GDP in 2018 = 2975

Growth rate in Real GDP = [(2975 - 1500) / 1500] * 100 = 98.33%

  • Nominal money in 2008 = 60

Nominal money in 2018 = 500

Growth rate in Nominal Money = [(500 - 60) / 60] * 100 = 733.33%

  • Nominal GDP = price * Real GDP

In 2008, Price = 60 / 1500 = 0.04

In 2008, Price = 500 / 2975 = 0.168

Thus Inflation rate = [(0.168 - 0.04) / 0.04] * 100 = 320%